[figure: 9.7 cm]

MAGNETIC PROPERTIES OF THORIUM FERRICYANIDE

R. Göbl, M. Zentková, V. Kavečanský, J. Kováč, A. Zentko Institute of Experimental Physics SAS, Watsonova 47, 043 53 Košice, Slovakia

In the last few years different groups have obtained encouraging results in the field of molecular based ferromagnets. One of the simplest and most convenient molecular building blocks can be found in hexacyanometallate family [B(CN)_{6}]^{n−}, where B is transition metal and n equals 3 and 4 respectively [1, 2, 3]. The aim of the present paper is to supply information on the magnetic and structural properties of Th_{3}[Fe(CN)_{6}]_{4}.15 H_{2}O compound. As Th^{4+} ion has no net moment, the system is very simple, with only one type of magnetic ion Fe^{3+}.

Literature

1. S. Ferlay, T. Mallah, R. Ouahes, P. Veillet, Nature, 378 (1995) 701

2. T. Mallah, S. Thiebaut, M. Verdaguer, P. Veillet, Science, 262 (1993) 1554

3. R.Klenze, B. Kanellakopulos, G. Trageser, H.H. Eysel, J.Chem.Phys. 72 (1980) 5819

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Mária Zentková

Address for correspondence : Institute of Experimental Physics SAS Watsonova 47, 043 53 Košice Slovakia

Email address : zentkova@saske.sk